# Online Collision Risk Estimation via Monocular Depth-Aware Object Detectors and Fuzzy Inference

### v1

@@ Abstract
This paper presents a monitoring framework that infers the level of autonomous vehicle (AV) collision risk based on its object detector’s performance using only monocular camera images. Essentially, the framework takes two sets of predictions produced by different algorithms and associates their inconsistencies with the collision risk via fuzzy inference. The first set of predictions is obtained through retrieving safety-critical 2.5D objects from a depth map, and the second set comes from the AV’s 3D object detector. We experimentally validate that, based on Intersection-over-Union (IoU) and a depth discrepancy measure, the inconsistencies between the two sets of predictions strongly correlate to the safety-related error of the 3D object detector against ground truths. This correlation allows us to construct a fuzzy inference system and map the inconsistency measures to an existing collision risk indicator. In particular, we apply various knowledge- and data-driven techniques and find using particle swarm optimization that learns general fuzzy rules gives the best mapping result. Lastly, we validate our monitor’s capability to produce relevant risk estimates with the large-scale nuScenes dataset and show it can safeguard an AV in closed-loop simulations.
This work has been submitted to the IEEE for possible publication. Copyright may be transferred without notice, after which this version may no longer be accessible.
@@ I Introduction
In the literature, in fact, one can find various run-time monitoring techniques. Focusing on planning and control, many work derive collision risks based on ego and traffic information, e.g., driving path deviation or time-to-collision to other agents [4]. These approaches, nonetheless, often assume perfect perception, which is generally not the case. In this work, we relax such an assumption and attempt to identify hazards as early as possible in an AV software stack, such that the controller can trigger a safety maneuver in time. Holding a similar mindset, several studies have suggested to monitor the object detection function. For instance, some propose algorithms that check the spatial or temporal consistency of the set of detected objects [5, 6]. Despite effective, one crucial limitation of the existing work is the relevance between the identified anomalies and the actual safety risk of the AV. For example, these monitors may raise a warning for an object that is far from the AV, which actually poses a low risk. Likewise, they only examine the set of detected objects and ignore potential misses (i.e., false negatives of the object detector), which are more likely risk-relevant.
@@
Figure 1: The overall monitoring framework using alignment measures between two sets of predictions to infer ego vehicle’s collision risks. The dashed lines mark an offline optimization process using a previously presented risk-correlating metric, USC [7].
@@
Hence, in this work, we aim to find safety-related errors of the object detector and use them to characterize an AV’s collision risks. Fig. 1 depicts the overall framework. In particular, our previous work presented a design-time safety-focused metric, called uncompromising spatial constraints (USC), which highly correlates with AV collision rates [7]. This work’s objective, therefore, is to reproduce it and generate a relevant risk level during operation. To achieve this, there are two challenges:
@@
  • The first and main challenge lies in the lack of ground-truth labels at run time. To overcome it, we ask the critical question whether employing a separate object retrieval pipeline and measuring the inconsistencies from the original object detector’s predictions can serve as a proxy to the ground truths. Specifically, considering cost factors and the recent breakthrough in monocular depth estimation, we employ the state-of-the-art ZoeDepth [8] and implement image processing techniques to retrieve safety-critical objects. Our key insight, thereby, is an experimental validation that confirms the inconsistencies between the two sets of predictions, measured by Intersection-over-Union (IoU) and depth discrepancy, are closely linked to safety-related errors in the 3D object detector when compared to ground-truth data.
@@
  • With the confirmed correlation, the second challenge is how to associate them with the desired risk quantifier, USC. Our solution is to use fuzzy logic, which can flexibly model non-linear functions while tolerating potential imprecision [9]. Concretely, we build a fuzzy inference system (FIS) using three distinct knowledge- and data-driven approaches and finally find the one adopting the global particle swarm optimization algorithm gives the best result within a separate testing dataset.
@@
To demonstrate the efficacy of our framework, we run it with a state-of-the-art 3D object detector, PGD [10], across the large-scale nuScenes dataset [11] in six conditions, including nominal scenes, night scenes, scenes with rare objects, and scenes augmented with rain, snow, or Gaussian noises. In addition, we implement our monitor on a baseline AV in simulation and showcase that it can indeed infer a useful risk indicator that helps protect the AV. Altogether, our work can be seen as a novel attempt to derive collision risk estimates from the object detection function using a single data source, i.e., images from one monocular camera. Moreover, as we shall show in later sections, the framework offers good interpretability and adaptability, allowing developers to continuously analyze and improve the monitor.
@@ II Related Work
Uncertainty quantification aims at estimating the confidence level of a model in its predictions. Beyond classical methods such as the softmax output [17], Monte Carlo dropout [18], or ensemble models [19], recent studies attempt to formulate uncertainty-aware loss functions and directly estimate the variance of the object detector’s predictions [20, 21]. Still, the main challenge in uncertainty quantification is its correlation with the actual model errors or system risks during operation. While research continues to investigate related uncertainty calibration techniques, we take a distinct yet analogous approach to infer system risks from model errors and achieve calibrated output via the adaptable FIS.
Similar in not explicitly modelling uncertainty, out-of-distribution (OoD) detection studies attempt to find other possibility to flag a data. For example, Du et al. models class-specific features with Gaussian distributions and generate OoD data to train an auxiliary OoD detector [22]. Recently, Wu et al. find it more effective to store features of training data into hyper-rectangles and use them as permissible ranges for new data [23]. However, it is hard to derive semantic information such as system risks from these approaches. Our usage of a depth estimator can naturally inform the monitor about such risks based on distance.
Accordingly, performance monitoring approaches checking object detector output lie the closest to our work. Existing approaches can be categorized into two groups. The first group makes use of learning-based models. As an example, Rahman et al. [24] train a neural network (NN) aside the original object detector to estimate the typical mean Average Precision (mAP) metric. Although they achieve great regression results, mAP might not be the suitable metric for safety or collision risk indication at each frame, as pointed out by a later study [25]. In addition, there are still concerns about NN robustness and interpretability. Our approach to regress a risk-relevant metric with fuzzy logic helps alleivate these concerns. Lastly, the second group of monitoring approaches uses formal methods and specifies rules such as the maintenance of class labels and the consistency of object locations across image frames [5, 6]. Their main constraint, as introduced, is the scope focusing on the set of predictions from an object detector only. This loses the possibility to check for detection misses. By contrast, our framework handles two sets of predictions, with one focusing on the safety-critical objects to check on the other.
Finally, our approach is related to data fusion using multiple sensing/perception channels. Readers may refer to a recent survey for more comprehensive understanding [26]. Compared to the mass literature, the crux of our work is the orientation towards risk indication with perception processes using only monocular camera images.
@@ III The Risk Estimation Framework
Figure 2: We implement a sequence of image processing techniques to find the safety-critical objects from a given depth map. The images from left to right depicts (a) the given depth map, (b) inverse of the depth map, (c) foreground removal on the inverse, and (d) bounding boxes rendered on the original image. The first three images have been normalized to the intensity range $[0,255]$ for visualization.
To iterate, we attempt to indicate relevant collision risks from object detection results during run time. In the following, Sec. III-A introduces how to attain 2.5D objects from a depth map and validates that comparing them with 3D object detection predictions gives valid proxies to ground-truth evaluation. Later, Sec. III-B presents the construction and optimization of the FIS that associate the comparison measures to the output risk level.
@@ III-A Depth-based object retrieval and consistency check
As mentioned, observing its potential to hint the whereabouts of objects, we employ a depth estimator that provides absolute depth maps [8]. Based on the depth maps, a typical object retrieval process consists of two main steps: (i) Apply the Canny edge detecting algorithm to mark the pixels that have large gradient magnitudes [27]; (ii) Find the objects by linking adjacent highlighted pixels into contours [28].
@@ III-A1 Preprocessing techniques
In our work, we exploit the depth information in driving contexts and implement the following two techniques before applying the typical process to facilitate better object retrieving results. First, we take the inverse of the raw depth map so as to obtain stronger gradients for pixels closer to the ego vehicle. To explain, with $I_{1}(x,y)$ and $I_{2}(x,y)$ being two intensity images linearly normalized from the raw depth map $D(x,y)$ and its inverse $\frac{1}{D(x,y)}$ respectively, their gradient magnitudes can be written with:
@@
One may notice that the foreground close to the ego vehicle is also highlighted after taking the inverse. Fortunately, as an AV always keeps a similar driving perspective, we can remove it by computing the average of all inverse depth map in the training data and subtract the average from each map. Together with the first technique, this results in a strong emphasis on the close objects, as shown in Fig. 2(c). Lastly, following the mentioned typical process, we can effectively retrieve the most safety-critical objects in a driving scene. Fig. 2(d) shows an example result with bounding boxes fitted.
@@ III-A2 Measuring alignment with object detector predictions
Using the retrieved safety-critical objects, we now compare them with the predictions of the underlying object detector. Specifically, we focus on the localization attributes and measure how spatially aligned the object detector’s predictions are with the safety-critical objects.
@@
More formally, we first represent the safety-critical objects with 2.5D information from the retrieved bounding boxes and the original depth map That is, we denote the set of $M$ safety-critical objects as $\tilde{\mathbf{S}}\mathrel{\stackrel{{\scriptstyle\makebox[0.0pt]{\mbox{\scriptsize def}}}}{{:=}}}\{\hat{\mathbf{S}}_{m}\,|\,m=1,\dots,M\}$, where each $\hat{\mathbf{S}}\mathrel{\stackrel{{\scriptstyle\makebox[0.0pt]{\mbox{\scriptsize def}}}}{{:=}}}(x_{\mathbf{S}},y_{\mathbf{S}},w_{\mathbf{S}},h_{\mathbf{S}},d_{\mathbf{S}})$, with $(x_{\mathbf{S}},y_{\mathbf{S}})$ being the center of the bounding box, $w_{\mathbf{S}}$ the width in the x-axis, $h_{\mathbf{S}}$ the height in the y-axis, and $d_{\mathbf{S}}$ the smallest depth value in the range of the bounding box.
@@
Correspondingly, we represent the set of $N$ predicted objects as $\tilde{\mathbf{P}}\mathrel{\stackrel{{\scriptstyle\makebox[0.0pt]{\mbox{\scriptsize def}}}}{{:=}}}\{\hat{\mathbf{P}}_{n}\,|\,n=1,\dots,N\}$, where each $\hat{\mathbf{P}}\mathrel{\stackrel{{\scriptstyle\makebox[0.0pt]{\mbox{\scriptsize def}}}}{{:=}}}(x_{\mathbf{P}},y_{\mathbf{P}},w_{\mathbf{P}},h_{\mathbf{P}},d_{\mathbf{P}})$. In practice, a 2D bounding box $\mathbf{P}$ without the depth attribute can be obtained from a 3D counterpart via perspective projection, and $d_{\mathbf{P}}$ can be concretized as the closest distance from the 3D bounding box to the AV’s position.
@@
Now, to measure the spatial alignment between the two sets $\tilde{\mathbf{P}}$ and $\tilde{\mathbf{S}}$, we consider two factors, namely the overlap of bounding boxes in the image plane and the depth discrepancy in the driving direction. Visualized in the side figure, these two factors have been shown highly related to safety in a recent analysis [29]. In practice, we first run the standard Hungarian algorithm to find the closest pairs in terms of their center distances and, for each pair, compute the commonly used Intersection-over-Union (IoU) as well as a simple relative depth discrepancy (RDD) measure:
@@
$\mathsf{RDD}(d_{\mathbf{P}},d_{\mathbf{S}})\mathrel{\stackrel{{\scriptstyle\makebox[0.0pt]{\mbox{\scriptsize def}}}}{{:=}}}\min(1,\frac{|d_{\mathbf{P}}-d_{\mathbf{S}}|}{d_{\mathbf{S}}})\in[0,1].$ (3)
@@
We additionally apply matching thresholds on the two measures, $\mathsf{IoU}>\alpha$ and $\mathsf{RDD}<\beta$, so as to avoid actually distant pairs. For any unmatched safety-critical objects, the worst scores are assigned, i.e., $\mathsf{IoU}=0$ and $\mathsf{RDD}=1$. Lastly, we aggregate the information in one frame by taking the average of the two measures for all safety-critical objects, denoted as $\widetilde{\mathsf{IoU}}$ and $\widetilde{\mathsf{RDD}}$.
@@ III-A3 Interim validation
To show the computed measures, $\widetilde{\mathsf{IoU}}$ and $\widetilde{\mathsf{RDD}}$, are potential for risk inference at the next stage, we now replace $\tilde{\mathbf{S}}$ with ground-truth annotations $\tilde{\mathbf{G}}$ in the above alignment measuring process and computes the correlation coefficient between the measures obtained with the safety-critical objects $\tilde{\mathbf{S}}$ and those with the ground truths $\tilde{\mathbf{G}}$.
For implementation, we use the nuScenes [11] dataset, the state-of-the-art PGD detector [10], as well as ZoeDepth estimator [8]. As our goal is to estimate relevant collision risks, we focus on objects that are situated within 20 meters from the ego vehicle.
TABLE I: Correlation between “measuring the alignment of the object detector’s predictions $\tilde{\mathbf{P}}$ and depth-retrieved objects $\tilde{\mathbf{S}}$” and “evaluating the predictions $\tilde{\mathbf{P}}$ with ground truths $\tilde{\mathbf{G}}$”. ($\alpha$, $\beta$) denotes the matching thresholds for $\mathsf{IoU}$ and $\mathsf{RDD}$, respectively.

### v2

# FuzzRisk: Online Collision Risk Estimation for Autonomous Vehicles based on Depth-Aware Object Detection via Fuzzy Inference

@@ Abstract
This paper presents a novel monitoring framework that infers the level of collision risk for autonomous vehicles (AVs) based on their object detection performance. The framework takes two sets of predictions from different algorithms and associates their inconsistencies with the collision risk via fuzzy inference. The first set of predictions is obtained by retrieving safety-critical 2.5D objects from a depth map, and the second set comes from the ordinary AV’s 3D object detector. We experimentally validate that, based on Intersection-over-Union (IoU) and a depth discrepancy measure, the inconsistencies between the two sets of predictions strongly correlate to the error of the 3D object detector against ground truths. This correlation allows us to construct a fuzzy inference system and map the inconsistency measures to an AV collision risk indicator. In particular, we optimize the fuzzy inference system towards an existing offline metric that matches AV collision rates well. Lastly, we validate our monitor’s capability to produce relevant risk estimates with the large-scale nuScenes dataset and demonstrate that it can safeguard an AV in closed-loop simulations.
© 2025 IEEE. Personal use of this material is permitted. Permission from IEEE must be obtained for all other uses, in any current or future media, including reprinting/republishing this material for advertising or promotional purposes, creating new collective works, for resale or redistribution to servers or lists, or reuse of any copyrighted component of this work in other works.
@@ I Introduction
In the literature, in fact, one can find various run-time AV monitoring techniques. Focusing on planning and control, many works draw collision risks based on ego and traffic information, e.g., deviation of the driving path or time to collision towards other agents [4]. However, these approaches assume perfect perception, which is usually not the case. Our work attempts to relax the assumption and identify hazards as early as possible at the perception phase of the AV software stack. Several studies have done so similarly, monitoring the object detection function, given its pivotal role for an AV to operate. Their algorithms, for example, check the spatial or temporal consistency of the set of detected objects [5, 6]. While their results appear effective, one crucial limitation is the relevance between the identified anomalies and the actual safety risk of the AV. For instance, these object detection monitors may raise a warning for an object that is actually far from the AV and less relevant. Likewise, they only examine the set of detected objects and ignore potential misses (i.e., false negatives of the object detector), which possibly pose greater risks to safety.
@@
Figure 1: Our overall monitoring framework uses alignment measures between two sets of predictions to infer ego vehicle’s collision risks. The dashed lines mark an offline optimization process using a previously presented risk-correlating metric, USC [7].
@@
Hence, in this work, we aim to find safety-relevant errors of the object detector and use them to characterize an AV’s collision risks. Fig. 1 depicts the overall framework. In particular, our previous work presented a design-time safety-oriented metric, uncompromising spatial constraints (USC), and showed that it matches AV collision rates well using simulations [7]. This work then utilizes USC during offline optimization to eventually obtain a reliable collision risk indicator that can help protect the AV in operation. To achieve this, there are two challenges:
@@
  • The first and main challenge lies in the lack of ground-truth labels at run time. To overcome it, we ask the critical question of whether employing a separate object retrieval pipeline and measuring the inconsistencies from the original object detector’s predictions can serve as a proxy to the ground truths. Considering cost factors and the recent breakthrough in monocular depth estimation, we employ the state-of-the-art ZoeDepth [8] and implement image processing techniques to retrieve safety-critical objects. Our key insight, thereby, is an experimental validation that confirms the inconsistencies between the two sets of predictions, measured by Intersection-over-Union (IoU) and depth discrepancy, are closely linked to the errors of the 3D object detector against ground truths.
@@
  • With the confirmed correlation, the second challenge is how to associate them with the desired risk quantifier. Our solution is to use fuzzy logic, which can flexibly model non-linear functions while tolerating potential imprecision [9]. More concretely, we build a fuzzy inference system (FIS) using three distinct knowledge- and data-driven approaches and finally find the global particle swarm optimization algorithm gives the best results.
@@
To demonstrate the efficacy of our framework, we run it with a state-of-the-art 3D object detector, PGD [10], across the large-scale nuScenes dataset [11] in six conditions, including nominal scenes, night scenes, scenes with rare objects, and scenes augmented with rain, snow, or Gaussian noises. In addition, we implement our monitor on a baseline AV in simulation and showcase that it can indeed infer a useful risk indicator that helps protect the AV. Altogether, our work can be seen as a novel attempt to derive safety-relevant collision risk estimates from the object detection function. Moreover, as we shall show in later sections, the framework offers good interpretability and adaptability, allowing developers to continuously analyze and improve the monitor.
@@ II Related Work
Uncertainty quantification aims to estimate a model’s confidence level in its predictions. Beyond classical methods such as the softmax output [17], Monte Carlo dropout [18], or ensemble models [19], recent studies attempt to formulate uncertainty-aware loss functions and directly estimate the variance of the object detector’s predictions [20, 21]. Still, the main challenge in uncertainty quantification is its correlation with the actual model errors or system risks during operation. While research continues to investigate related uncertainty calibration techniques, we take an analogous yet distinct approach to infer system risks from model errors and calibrate the output via an adaptable FIS.
Out-of-distribution (OoD) detection studies seek solutions other than model confidence estimation to identify anomalous data. For example, Du et al. models class-specific features with Gaussian distributions and generates samples to train an auxiliary OoD detector [22]. Recently, Wu et al. find it more effective to store features of training data into hyper-rectangles and use them as permissible ranges to monitor inputs [23]. However, it is hard to derive semantic information such as system risks from these feature-based approaches. Our usage of a depth estimator can naturally signal the monitor about such risks based on distance information.
Accordingly, performance monitors that operate on the object detector’s outputs lie the closest to our work. Existing approaches can be categorized into two groups. The first group makes use of learning-based models. As an example, Rahman et al. [24] train a neural network (NN) aside from the original object detector to estimate the typical mean Average Precision (mAP) metric. Although they achieve great regression results, mAP might not be the suitable metric for safety or collision risk indication at each frame, as pointed out by a later study [25]. In addition, there are still concerns about NN robustness and interpretability. Our proposal to regress a risk-relevant metric with fuzzy logic helps alleviate these concerns. Lastly, the second group of monitoring approaches uses formal methods and specifies rules such as the consistency of predicted class labels or object locations across image frames [5, 6]. Their main constraint, as introduced, is the scope focusing on the set of predictions from an object detector only. This loses the possibility to check for detection misses. By contrast, our framework handles two sets of predictions, with one focusing on the safety-critical objects to check on the other.
Finally, our approach is related to data fusion using multiple sensing/perception channels. Readers may refer to a recent survey for a more comprehensive understanding [26]. Compared to the mass literature, the crux of our work is the orientation towards risk indication with perception processes using only camera images.
@@ III The Risk Estimation Framework
Figure 2: We implement a sequence of image processing techniques to find the safety-critical objects from a given depth map. The images from left to right depict (a) the given depth map, (b) the inverse of the depth map, (c) foreground removal on the inverse, and (d) the bounding boxes rendered on the original image. The first three images have been normalized to the intensity range $[0,255]$ for visualization.
To iterate, we attempt to indicate relevant collision risks from object detection results during run time. In the following, Sec. III-A introduces how to attain 2.5D objects from a depth map and validates that comparing them with 3D object detection predictions constitutes a valid proxy to ground-truth evaluation. Later, Sec. III-B presents the construction and optimization of the FIS that associates the comparison measures to the output risk level.
@@ III-A Depth-based object retrieval and consistency check
As introduced, observing its potential to hint at the whereabouts of objects, we employ a depth estimator that provides absolute depth maps [8]. Based on the depth maps, a typical object retrieval process consists of two main steps: (i) Apply the Canny edge detecting algorithm to mark the pixels that have large gradient magnitudes [27]; (ii) Find the objects by linking adjacent highlighted pixels into contours [28].
@@ III-A1 Preprocessing techniques
In our work, we exploit the depth information in driving contexts and implement the following two techniques before applying the aforementioned typical process to facilitate better object retrieving results. First, we take the inverse of the raw depth map so as to obtain stronger gradients for pixels closer to the ego vehicle. To explain, with $I_{1}(x,y)$ and $I_{2}(x,y)$ being two intensity images linearly normalized from the raw depth map $D(x,y)$ and its inverse $\frac{1}{D(x,y)}$ respectively, their gradient magnitudes can be written with:
@@
One may notice that the foreground close to the ego vehicle is also highlighted after taking the inverse. Fortunately, as an AV always keeps a similar driving perspective, we can remove it by computing the average of all inverse depth maps in the training data and subtracting the average from each map. Together with the first technique, this results in a strong emphasis on the close objects, as shown in Fig. 2(c). Lastly, following the typical process (i.e., Canny edge detection and contour linking), we can effectively retrieve the most safety-critical objects in a driving scene. Fig. 2(d) shows an example with fitted bounding boxes.
@@ III-A2 Measuring alignment with object detector predictions
Using the retrieved safety-critical objects, we now compare them with the predictions of the underlying object detector. In particular, we focus on the localization attributes and measure how spatially aligned the object detector’s predictions are with the safety-critical objects.
@@
To start, we represent the safety-critical objects with 2.5D information from the retrieved bounding boxes, e.g., Fig. 2(d), and the original depth map, e.g., Fig. 2(a). More formally, we denote the set of $M$ safety-critical objects in one frame as $\tilde{\mathbf{S}}\mathrel{\stackrel{{\scriptstyle\makebox[0.0pt]{\mbox{\scriptsize def}}}}{{:=}}}\{\hat{\mathbf{S}}_{m}\,|\,m=1,\dots,M\}$, where each $\hat{\mathbf{S}}\mathrel{\stackrel{{\scriptstyle\makebox[0.0pt]{\mbox{\scriptsize def}}}}{{:=}}}(x_{\mathbf{S}},y_{\mathbf{S}},w_{\mathbf{S}},h_{\mathbf{S}},d_{\mathbf{S}})$, with $(x_{\mathbf{S}},y_{\mathbf{S}})$ being the center of the bounding box, $w_{\mathbf{S}}$ the width in the x-axis, $h_{\mathbf{S}}$ the height in the y-axis, and $d_{\mathbf{S}}$ the smallest depth value in the area of the bounding box.
@@
Correspondingly, we represent the set of $N$ predicted objects in the frame as $\tilde{\mathbf{P}}\mathrel{\stackrel{{\scriptstyle\makebox[0.0pt]{\mbox{\scriptsize def}}}}{{:=}}}\{\hat{\mathbf{P}}_{n}\,|\,n=1,\dots,N\}$, where each $\hat{\mathbf{P}}\mathrel{\stackrel{{\scriptstyle\makebox[0.0pt]{\mbox{\scriptsize def}}}}{{:=}}}(x_{\mathbf{P}},y_{\mathbf{P}},w_{\mathbf{P}},h_{\mathbf{P}},d_{\mathbf{P}})$. In practice, a 2D bounding box $\mathbf{P}$ without the depth attribute can be obtained from a 3D counterpart via perspective projection, and $d_{\mathbf{P}}$ can be concretized as the closest distance from the 3D bounding box to the AV’s position.
@@
Now, to measure the spatial alignment between the two sets $\tilde{\mathbf{P}}$ and $\tilde{\mathbf{S}}$, we consider two factors, namely the overlap of bounding boxes in the image plane and the depth discrepancy in the driving direction. Visualized in the side figure, these two factors have been shown to be highly relevant to safety in a recent analysis [29]. In practice, we first run the standard Hungarian algorithm to find the closest pairs in terms of their center distances and, for each pair, compute the commonly used Intersection-over-Union (IoU) as well as a simple relative depth discrepancy (RDD) measure:
@@
$\mathsf{RDD}(d_{\mathbf{P}},d_{\mathbf{S}})\mathrel{\stackrel{{\scriptstyle\makebox[0.0pt]{\mbox{\scriptsize def}}}}{{:=}}}\min(1,\frac{d_{\mathbf{P}}-d_{\mathbf{S}}}{d_{\mathbf{S}}})\in[-1,1].$ (3)
@@
We additionally apply matching thresholds on the two measures, $\mathsf{IoU}>\alpha$ and $|\mathsf{RDD}|<\beta$, so as to avoid actually distant pairs. For any unmatched safety-critical objects, the worst scores are assigned, i.e., $\mathsf{IoU}=0$ and $\mathsf{RDD}=1$. Lastly, we aggregate the information in one frame by taking the average of the two measures for all safety-critical objects, denoted as $\widetilde{\mathsf{IoU}}$ and $\widetilde{\mathsf{RDD}}$.
@@ III-A3 Interim validation
To show the computed measures, $\widetilde{\mathsf{IoU}}$ and $\widetilde{\mathsf{RDD}}$, are potential for risk inference at the next stage, we now replace $\tilde{\mathbf{S}}$ with ground-truth annotations $\tilde{\mathbf{G}}$ in the above alignment measuring process and computes the Pearson correlation coefficient between the measures obtained with the safety-critical objects $\tilde{\mathbf{S}}$ and those with the ground truths $\tilde{\mathbf{G}}$.
For implementation, we use the nuScenes [11] dataset, the state-of-the-art PGD detector [10], as well as ZoeDepth estimator [8]. As the most prominent collision risks stem from nearby objects, we focus on those situated within 20 meters of the ego vehicle.
TABLE I: The positive correlation between “measuring the alignment of the object detector’s predictions $\tilde{\mathbf{P}}$ and depth-retrieved objects $\tilde{\mathbf{S}}$” and “evaluating the predictions $\tilde{\mathbf{P}}$ with ground truths $\tilde{\mathbf{G}}$.” The correlation increases as the matching thresholds are relaxed in the second row. ($\alpha$, $\beta$) denotes the matching thresholds for $\mathsf{IoU}$ and $\mathsf{RDD}$, respectively.